### v4

# Beyond Single-Device Constraints: A System-Level Theoretical Framework for High-Performance Single-Photon Detection at Room Temperature

Abstract
Photon detection, a fundamental quantum technology, is traditionally treated as a fixed device-level operation governed by intrinsic properties of single-photon detectors (SPDs). High-performance detection has therefore largely relied on superconducting technologies whose requirement for cryogenic operation imposes substantial infrastructure constraints, limiting scalable deployments. Here, the enhanced single-photon detection (ESPD) framework is presented as a system-level theoretical paradigm that shifts photon detection from device-centric optimization to an integrated quantum-information-processing task, by reformulating it as an iteratively enhanced process integrating state-preparation, controlled operations, projective measurements, and multi-copy analysis. ESPD enables systematic performance enhancement through architectural design rather than material modification, thereby circumventing superconducting components, allowing high-performance detection using exclusively room-temperature hardware. Numerical simulations grounded in physically motivated parameters indicate that the ESPD framework can upgrade a conventional room-temperature SPD to effective DE exceeding 93% and DCR below $10^{-9}$, which are comparable to state-of-the-art superconducting SPDs and can significantly relax the minimal tolerable channel transmission rate in quantum communications. While physical realization would require further component integration, this work establishes a rigorous theoretical framework for enhancing detection performance through architectural quantum-information principles, providing a general blueprint for transcending device-level constraints and guiding the development of next-generation room-temperature quantum technologies.
keywords:
Single-photon detection, System-level quantum architecture, Iterative quantum information processing, Room-temperature quantum technologies, Photonic integration
Hao Shu*
[…]
{affiliations}
Dr. Hao Shu
Address: Sun Yat-Sen University, Guangzhou, China
Email Address: Hao_B_Shu@163.com (Primary), shuh6@mail.sysu.edu.cn (Institutional)
1 Introduction
Single-photon detection plays a central role in quantum optics and serves as the bedrock for a wide range of quantum-information applications. Conventionally, the detection process is treated as a fixed device-level operation primarily determined by the material properties of single-photon detectors (SPDs). The primary performance metrics characterizing an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$. The former directly dictates the success probability of quantum operations conditioned on detection events, while the latter imposes a fundamental signal-to-noise constraint in high-attenuation scenarios.
In terms of DE, high-performance detection is indispensable for retrieving meaningful outcomes in quantum experiments. In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the overall success probability decreases exponentially with imperfect DE [1, 2, 3]. In fundamental tests of quantum nonlocality, verifying violations of the Clauser-Horne-Shimony-Holt or Eberhard inequalities requires DE exceeding strict thresholds to close detection loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12]. In heralded single-photon sources, where one photon of an entangled pair signals the presence of the other, the achievable heralding rate is also strictly limited by DE [13, 14, 15, 16].
With respect to DCR, a particularly stringent constraint arises in quantum communication, notably in quantum key distribution (QKD). The security of practical QKD protocols relies on maintaining the quantum bit error rate (QBER) below protocol-dependent thresholds that are significantly lower than 50% [17, 18, 19, 20]. However, optical communication channels inevitably induce exponential photon attenuation with distance [21, 22, 23, 24]. When the effective signal rate, determined by channel transmission for a given DE, approaches the DCR floor of the SPDs, detection events become dominated by dark counts, which contribute a 50% error rate, causing the QBER to rise precipitously. Consequently, the achievable secure distance of QKD systems is ultimately bounded by the DCR of the employed SPDs [25, 26, 27].
Motivated by these critical demands, the development of high-performance SPDs has remained a primary research focus. While conventional semiconductor-based SPDs offer the advantage of non-cryonic operation, they typically suffer from an intrinsic trade-off between limited DE and relatively high DCR. For instance, a frequency upconversion SPD demonstrated in 2004 operated at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP SPD reported in 2017 operated at 223 K achieved a DE of about 27.5% with a DCR of $10^{-6}$ [29]. Such specifications often fall short of the requirements for advanced quantum protocols, which frequently demand DE exceeding 90% and DCR below $10^{-7}$.
To meet these stringent requirements, substantial effort has shifted toward designing SPDs with superconducting technologies, most notably superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These devices have successfully pushed DE beyond 90% while suppressing DCR by several orders of magnitude compared to semiconductor counterparts, thereby enabling demanding applications such as long-distance QKD [19, 34, 35, 36]. Recent representative benchmarks of state-of-the-art superconducting SPDs are summarized in Table 1.
Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs.
Year Researchers (et al.) Operation Temperature Detection Efficiency Dark Count Rate
2020 P. Hu   [30] 2.10 K 95.0 % $0.5\times 10^{-5}$
2021 J. Chang [31] 2.50 K 99.5 % $1.1\times 10^{-3}$
2021 G. Z. Xu [32] 0.84 K 92.2 % $3.6\times 10^{-5}$
2023 I. Craiciu [33] 0.90 K 78.0 % $1.0\times 10^{-7}$
2024 I. Charaev [37] 20.0 K 7.60 % $1.0\times 10^{-3}$
However, the superior performance of superconducting SPDs comes at the cost of cryogenic operation. This imposes stringent environmental requirements and necessitates bulky, power-intensive cooling infrastructure. The significant degradation of performance at elevated temperatures highlights a strong trade-off between detection performance and operating conditions. These constraints present major barriers to scalable deployment, particularly in field applications, satellite payloads, and scenarios with strict size, weight, and power limitations where cryogenic infrastructure is impractical.
To address the limitations of room-temperature SPDs without resorting to cryogenic conditions, the enhanced single-photon detection (ESPD) framework is introduced, shifting the focus from device-centric optimization to a system-level architectural paradigm. Within this framework, single-photon detection is reformulated as a composite quantum-information-processing task that integrates state preparation, controlled quantum operations, projective measurements, and multi-copy decision analysis, rather than relying solely on the material properties of a single SPD. Through structured system-level integration, the ESPD framework enables effective enhancement of detection performance using exclusively room-temperature components.
Numerical simulations, grounded in physically motivated parameters drawn from commercially available devices and experimentally reported SPDs, indicate that, under an exclusively room-temperature component setting, the ESPD framework enables effective DE exceeding 93% together with effective DCR below $10^{-9}$, starting from a legacy SPD (for example, the one reported two decades ago with about 59% DE and $10^{-2}$ DCR [28]). These effective performance metrics are comparable to those of state-of-the-art superconducting SPDs and lead to a substantial relaxation of the minimal tolerable channel transmission rate in QKD.
While physical implementation would require further system-level integration of quantum technologies, the ESPD framework provides a structured theoretical basis for achieving high-performance photon detection at room temperature. By explicitly analyzing technological feasibility and the trade-off between resource overhead and operating temperature, the ESPD framework offers a feasible technological roadmap and practical guidance for future experimental efforts. It serves as a theoretical blueprint for overcoming intrinsic device limitations, with the potential to unlock a wide range of quantum applications in environmentally constrained scenarios.
2 Method and Result
This section details the architectural structure of the ESPD framework and formulates the corresponding detection process. The ESPD paradigm is introduced in Section 2.1, with theoretical formulations established in Section 2.2. Analytical approximations are developed in Section 2.3, and numerical simulations are presented in Section 2.4. Finally, implications for QKD are discussed in Section 2.5.
2.1 Paradigm of the Enhanced Single-Photon Detection
The ESPD architecture is formulated as a level-by-level framework in which the effective detection performance of a given SPD is enhanced through successive processing stages. Here, the enhancement is quantified in terms of DE and DCR, defined respectively as the probability of generating a positive detection event given a non-vacuum (single-photon) input and the probability of a positive report given a vacuum input.
Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with effective DE $\eta_{s}$ and effective DCR $d_{s}$. The iteration begins with an existing SPD, denoted as $\text{ESPD}_{0}$, characterized by the initial values $\eta_{0}=\eta$ and $d_{0}=d$.
With qubit logic, the schematic of the ESPD paradigm is illustrated in Fig. 1.
Figure 1: Schematic representation of the recursive ESPD architecture. The $(s+1)$-th level detector ($\text{ESPD}_{s+1}$) is constructed by integrating multiple copies of the $s$-th level detector ($\text{ESPD}_{s}$) with controlled gates. Here, $C(F_{i})$ are controlled gates, SP are state-preparation operations, measurements are projective measurements via basis $\{|0\rangle,|1\rangle\}$, and filters represent that the $|0\rangle$ path is aborted.
The construction of the $(s+1)$-th level detector, $\text{ESPD}_{s+1}$, is defined as follows. As a representative example, an implementation may be considered in which the degree of freedom (DOF) is photon polarization and the controlled operations correspond to controlled-NOT (C-NOT) gates.
  1. 1.
    DOF Selection: A manipulable DOF of the incoming photon is chosen (for example, polarization). The computational basis for this DOF is denoted by $\{|0\rangle,|1\rangle\}$. The signal is modeled as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle$ and $|1\rangle$ denote orthonormal states of a non-vacuum photon [^1] If the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents a non-vacuum state, reducing the description to qubit. In practical architectures, an internal DOF such as polarization, distinct from photon existence, is often more convenient..
  2. 2.
    State Preparation: A state-preparation operation is applied to initialize the selected DOF to $|1\rangle$. As a result, a non-vacuum signal is mapped to $|1\rangle$, while a vacuum input remains in state $|\emptyset\rangle$.
  3. 3.
    Controlled Operations: A set of $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates), are applied to couple the signal path with $n_{s+1}$ auxiliary photons, each initialized in state $|0\rangle$.
  4. 4.
    Auxiliary Measurement: Projective measurements are performed on the auxiliary outputs in basis $\{|0\rangle,|1\rangle\}$. Outcome paths corresponding to $|0\rangle$ are discarded, and only the $|1\rangle$ outcomes are retained for subsequent detection.
  5. 5.
    Detection: Detection is carried out on all remaining paths, including both the signal path and the auxiliary paths, using $s$-th level detectors ($\text{ESPD}_{s}$).
  6. 6.
    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detection events are registered among the $n_{s+1}+1$ detectors, consisting of $n_{s+1}$ auxiliary detectors and one detector on the signal path. Otherwise, a negative report is returned. Here, $k_{s+1}$ is a preselected threshold integer, and $n_{s+1}$ denotes the number of effective controlled operations [^2] The main discussions of the article assume deterministic controlled gates, for which $n_{s+1}$ equals the number of implemented operations. For probabilistic gates, $n_{s+1}$ refers to the number of successful operations after post-selection. In this case, the decision logic is naturally reformulated in terms of the fraction of positive detection outcomes among the effective detections exceeding a predefined threshold..
Physical intuition: From a detection perspective, each level of the ESPD framework increases the number of effective detection opportunities available to an input signal. In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level produces $n_{s+1}$ effective signal instances through the controlled modules, which are independently detected by $n_{s+1}$ SPDs. A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered. Under this idealized assumption, the effective DE increases with decreasing threshold $k_{s+1}$; for instance, when $k_{s+1}=1$, the effective DE is given by $1-(1-\eta_{s})^{n_{s+1}}$. Furthermore, the effective DCR is suppressed, since a false positive outcome requires dark-count events to occur in at least $k_{s+1}$ detectors, leading to an approximate scaling $d_{s}^{k_{s+1}}$. This physical intuition extends beyond the idealized limit when imperfections in the auxiliary components introduce errors that remain small compared to the activation probability of a genuine input signal. A rigorous formulation of this behavior, accounting for non-ideal devices and error propagation, is developed in Section 2.2.
2.2 Theoretical Formulations
The rigorous theoretical formulation of the ESPD framework is developed in this subsection. The relevant notations and assumptions are introduced first, followed by the exact analytical expressions and the approximation analysis.
2.2.1 Notations and Basic Assumptions
The parameters governing the system behavior are defined as follows.
As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively. Although $\eta_{s}$ may depend on the photon number distribution, it is approximated as a unified value for the given system for simplicity. The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled operations and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
For convenience, all controlled operations $C(F_{i})$ are assumed to be identical and are denoted simply as $C(F)$. They are implemented in a feedforward controlled configuration, such that $n_{s+1}$ represents both the number of effective modules and the total number of physical modules required at level $s+1$. However, the generalized ESPD paradigm remains valid for post-selection controlled gates (see Section 3.1.2 for a brief discussion). Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$). Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared). Let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively. Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection). The parameters $p$, $P$, and $Q$ are approximately assumed to be invariant across all $C(F)$ module operations.
It is crucial to note that $p$, $P$, and $Q$ represent the intrinsic performance metrics of a single, non-cascaded controlled module. These parameters are defined at the component level and remain invariant with respect to the number of cascaded modules. The cumulative effects of loss and error due to the system’s cascaded architecture will be fully incorporated into the derived system equations, such as $\eta_{s+1}$ in Eq.(6).
2.2.2 Parameter Approximation Assumptions
The assumptions below are employed solely to derive analytical approximations in Section 2.3, as well as the approximation part in Eq. (1) to (3). They are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations presented in this work. Although these assumptions entail relatively stringent requirements, they are physically motivated by realistic components and will be discussed in detail in Section 3.1.
In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$. Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds. It is further assumed that $d_{s}\ll P\eta_{s}$: If this condition were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant. Additionally, the transmission rate of a single controlled module is assumed to satisfy $p\approx 1$, and the SPAM error of the auxiliary path is assumed to satisfy $Q\ll 1$, consistent with the maturity of commercial optical components (see Section 3.1). Finally, considering practical implementation costs, $n_{s}$ and $k_{s}$ are taken to be moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
2.2.3 Formulations
Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum and vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector. They can be expressed as:
$P_{s+1}=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$ (1)
$Q_{s+1}$ $=Q[\eta_{s}+(1-\eta_{s})d_{s}]+(1-Q)d_{s}=Q\eta_{s}(1-d_{s})+d_{s}\approx Q\eta_{s}+d_{s}\leq Q+d_{s}$ (2)
Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum and vacuum output from the final controlled gate, respectively, then:
$P^{\prime}_{s+1}=\eta_{s}+(1-\eta_{s})d_{s}\approx\eta_{s},\ Q^{\prime}_{s+1}=d_{s}$ (3)
The DE and DCR for the $(s+1)$-th level are derived as follows.
For the DE calculation, a non-vacuum input is assumed, and the probability of obtaining at least $k_{s+1}$ positive reports out of $n_{s+1}+1$ total detections is evaluated. If the photon is lost exactly after the $i$-th $C(F)$, the probability of a final positive report is:
$P_{s+1,k_{s+1},i}=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$ (4)
$+Q^{\prime}_{s+1}\sum_{j_{1}+j_{2}=k_{s+1}-1}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
where the first and second terms correspond to scenarios that the signal path detector reports negative and positive, respectively.
If the photon survives all $C(F)$ operations, the probability of a final positive report is:
$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
By combining the above contributions, the DE of $\text{ESPD}_{s+1}$ is given by:
$\eta_{s+1}=p^{n_{s+1}}P_{s+1,k_{s+1}}+\sum_{i=1}^{n_{s+1}}p^{i-1}(1-p)P_{s+1,k_{s+1},i}$ (6)
For the DCR calculation, a vacuum input is assumed. The probability of a false positive is given by:
$d_{s+1}=$ $(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
Collectively, the DE and DCR evolution is governed by the non-linear time-variant dynamic system:
$(\eta_{s+1},d_{s+1})=G_{s+1}(\eta_{s},d_{s}),\ (\eta_{0},d_{0})=(\eta,d)$ (8)
where the map $G_{s+1}$ is defined by Eq. (1) to (7).
2.3 Analytical Approximations and Stability Analysis
While Eq. (8) describes the exact dynamics, an analytical solution is generally intractable. However, by employing the approximations in Eq. (1) to (3) together with the assumptions in Section 2.2.2, simplified bounds can be derived to elucidate the system behavior.
The evolution of the DCR is considered first. Eq. (7) can be reformulated as:
$d_{s+1}=$ $Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1-Q_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (9)
Proposition 1.
Define the function
$f(x)=a\binom{n}{k-1}x^{k-1}(1-x)^{n-k+1}+\sum_{j\geq k}\binom{n}{j}x^{j}(1-x)^{n-j}$ (10)
for $a\geq 0$, then $f(x)$ is monotonically increasing for $x\in[0,\frac{k-1}{n}]$.
Proof: For $x\in[0,\frac{k-1}{n}]$:
$\frac{d}{dx}f(x)=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$ (11)
$+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}(k-1-nx)+\sum_{j\geq k}\binom{n}{j}x^{j+1-k}(1-x)^{n-j-1}(j-nx)]$
$\geq$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}+\sum_{j\geq k}\binom{n}{j}x^{j+1-k}(1-x)^{n-j-1}](k-1-nx)\geq 0$
$\blacksquare$
By applying the upper bound from Eq. (2), substituting Eq. (3) into Eq. (9), and assuming $k-1\geq nx$, the following bound is obtained:
$d_{s+1}\leq$ $d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$ (12)
which is notably independent of $\eta_{s}$. Given $Q_{s},d_{s}\ll 1$ and moderate values of $n_{s+1}$ and $k_{s+1}$, it is assumed that $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$. This yields the approximation:
$d_{s+1}\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$ (13)
This provides an estimation for the DCR at level $s+1$.
Similarly, the DE can be lower-bounded by:
$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$ (14)
$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
which is independent of $d_{s}$. Therefore, the condition for DE enhancement at the subsequent level is determined by solving the inequality
$p^{n}(Px)^{k-1}[x\binom{n}{k-1}(1-Px)^{n-k+1}+\sum_{j\geq k}\binom{n}{j}(Px)^{j-k+1}(1-Px)^{n-j}]-x>0$ (15)
If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of $\text{ESPD}_{s}$ will converge to a root of the left-hand side of Eq. (15).
2.4 Numerical Simulations
While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design. In this section, numerical simulations of the ESPD paradigm are presented, indicating that substantial performance gains can be achieved even with moderate initial hardware specifications. For example, a three-level ESPD can upgrade a semiconductor SPD to effective performance metrics with DE $93\%$ and DCR $10^{-9}$, comparable to those of state-of-the-art SNSPDs without reliance on superconducting materials.
Simulation Parameters: The performance of the ESPD scheme is evaluated using two distinct baseline SPDs. The first represents an older one, characterized by parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29]. Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation[43, 38] (see Section 3.1 for parameter discussions). Supplementary simulations with degraded parameter settings are also performed to evaluate robustness in Section 3.1.
Performance Outcomes: Numerical performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig. 2. The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs. Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that are comparable to those of advanced superconducting SPDs.
Table 2: Performance of an ESPD initialized with a SPD from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
Para 1 Para 2 Para 3
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (4,1) $(97.4\%,\ 5.3\times 10^{-2})$ (6,1) $(98.4\%,\ 7.5\times 10^{-2})$ (8,1) $(98.6\%,\ 9.5\times 10^{-2})$
2 (4,2) $(98.2\%,\ 2.7\times 10^{-2})$ (6,3) $(96.6\%,\ 1.2\times 10^{-2})$ (8,4) $(95.1\%,\ 7.4\times 10^{-3})$
3 (4,2) $(98.0\%,\ 7.7\times 10^{-3})$ (6,3) $(95.8\%,\ 8.9\times 10^{-5})$ (8,4) $(93.6\%,\ 8.1\times 10^{-7})$
4 (4,2) $(97.9\%,\ 8.4\times 10^{-4})$ (6,3) $(95.6\%,\ 1.7\times 10^{-7})$ (8,4) $(93.4\%,\ 8.6\times 10^{-10})$
5 (4,2) $(97.8\%,\ 5.6\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
6 (4,2) $(97.8\%,\ 2.5\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
7 (4,2) $(97.8\%\ ,2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
Table 3: Performance of an ESPD initialized with a SPD from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
Para 1 Para 2 Para 3
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (4,1) $(76.9\%,\ 2.2\times 10^{-3})$ (6,1) (85.5%, $5\%,\ 3.4\times 10^{-3}) (8$,1) $(90.0\%,\ 4.5\times 10^{-3})$
2 (4,2) $(95.3\%,\ 1.2\times 10^{-4})$ (6,2) $(94.8\%,\ 3.8\times 10^{-4})$ (8,4) $(93.1\%,\ 1.7\times 10^{-7})$
3 (4,2) $(97.7\%,\ 2.6\times 10^{-5})$ (6,3) $(95.5\%,\ 1.4\times 10^{-7})$ (8,4) $(93.3\%,\ 8.4\times 10^{-10})$
4 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.4\times 10^{-10})$
5 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
6 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
7 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
(d) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
Figure 2: Performance of ESPD initialized with SPDs from [28] ((a), (b)) and [29] ((c), (d)), respectively.
2.5 Implications for QKD
As a representative application scenario, the implications of the ESPD framework for QKD are analyzed. For a QKD protocol characterized by a secure error threshold $e_{th}$ and a total error rate $e_{C}$ excluding detector contributions, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by[27]:
$\gamma=\frac{(1-2e_{th})d}{\eta[e_{th}-e_{C}+d(1-2e)]}\approx\frac{1-2e_{th}}{e_{th}-e_{C}}\frac{d}{\eta}$ (16)
Since the ESPD paradigm can substantially suppress the effective DCR $d$, the minimal tolerable channel transmission rate $\gamma$ is proportionately reduced. Furthermore, the simultaneous enhancement of the DE $\eta$ not only reduces $\gamma$ but also directly improves the achievable key rate at a fixed channel transmission rate.
Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD framework achieves this improvement without introducing additional sifting procedures that reduce the effective sifting efficiency. Here, QKD performance can be enhanced through the simultaneous suppression of the effective DCR and enhancement of the effective DE, without the need for auxiliary sifting operations. In this sense, the ESPD paradigm provides a more efficient system-level approach for improving QKD performance under practical constraints.
3 Discussion
In the absence of an experimental demonstration, we focus on the technological feasibility and practical robustness of the ESPD framework, and discuss its limitations as well as directions for future work.
3.1 Technological Feasibility and Performance under Degraded Parameters
The most significant advantage of the ESPD paradigm lies in its reliance on accessible, room-temperature components, effectively bypassing the cryogenic requirements of superconducting SPDs. The device requirements are discussed as follows.
3.1.1 State Preparations and Measurements
The flexibility of the ESPD architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one. In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be achieved at extremely low levels. Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
3.1.2 Controlled Gates
The primary technological prerequisites of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$.
Table 4: Performance of an ESPD initialized with a SPD from [28], with $p$ varied from 0.80 to 0.96. Simulations are performed for $(n,k)=(5,2)$ at each level, and all other settings remain consistent with those described in Section 2.4.
pp=0.80 pp=0.84 pp=0.88 pp=0.92 pp=0.96
Level (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$))
0 $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$
1 $(60.8\%,1.8\times 10^{-3})$ $(66.7\%,1.8\times 10^{-3})$ $(72.9\%,1.8\times 10^{-3})$ $(79.7\%,1.8\times 10^{-3})$ $(87.0\%,1.8\times 10^{-3})$
2 $(61.0\%,1.1\times 10^{-4})$ $(70.6\%,1.2\times 10^{-4})$ $(79.7\%,1.3\times 10^{-4})$ $(87.8\%,1.4\times 10^{-4})$ $(94.6\%,1.5\times 10^{-4})$
3 $(61.0\%,1.8\times 10^{-5})$ $(72.6\%,2.4\times 10^{-5})$ $(82.1\%,3.1\times 10^{-5})$ $(89.5\%,3.7\times 10^{-5})$ $(95.4\%,4.3\times 10^{-5})$
4 $(60.9\%,1.5\times 10^{-5})$ $(73.5\%,2.2\times 10^{-5})$ $(82.8\%,2.8\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
5 $(60.8\%,1.5\times 10^{-5})$ $(74.0\%,2.2\times 10^{-5})$ $(83.1\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
6 $(60.8\%,1.5\times 10^{-5})$ $(74.2\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
7 $(60.8\%,1.5\times 10^{-5})$ $(74.3\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
8 $(60.8\%,1.5\times 10^{-5})$ $(74.3\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
(a) DE $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
(b) DCR $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
Figure 3: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section 2.4.
A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity. It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module and, by definition, does not decrease with an increasing number of cascaded modules or ESPD levels. The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations. While the ESPD paradigm remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded. Specifically, if the post-selection efficiency of the controlled gate is $\frac{1}{N}$, the effective average transmission rate from one module to the next is reduced to $p^{N}$. Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate. Table 4 and Fig.3 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.). This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity should account for the correct operation rate of both the control and controlled paths. Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38]. Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% at non-cryonic components[39, 40, 41, 42]. However, it is also worth testing the effectiveness of the paradigm under degraded parameters to indicate its suitability for standard fabrication processes with lower technology requirements. Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5), and $P=0.40$ (see Table 6), with visual representation in Fig.4. All other settings remain consistent with those described in Section 2.4. The numerical results indicate the robustness of the scheme: Comparison of Table 5 to Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$. This robustness demonstrates that the technological requirements of the ESPD method are fully compatible with existing room-temperature quantum technologies.
Table 5: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.80$.
Para 1 Para 2 Para 3
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (4,2) $(77.8\%,\ 1.2\times 10^{-3})$ (6,2) $(88.5\%,\ 2.4\times 10^{-3})$ (8,2) $(92.3\%,\ 4.2\times 10^{-3})$
2 (4,2) $(91.3\%,\ 5.7\times 10^{-5})$ (6,3) $(92.4\%,\ 2.1\times 10^{-6})$ (8,4) $(91.0\%,\ 1.4\times 10^{-7})$
3 (4,2) $(95.7\%,\ 2.2\times 10^{-5})$ (6,3) $(93.3\%,\ 1.3\times 10^{-7})$ (8,4) $(90.6\%,\ 7.6\times 10^{-10})$
4 (4,2) $(96.4\%,\ 2.3\times 10^{-5})$ (6,3) $(93.5\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
5 (4,2) $(96.5\%,\ 5.5\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
6 (4,2) $(96.5\%,\ 2.4\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
7 (4,2) $(96.5\%,\ 2.3\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
8 (4,2) $(96.5\%,\ 2.3\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
Table 6: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under a low controlled gate fidelity, $P=0.40$.
Para 1 Para 2 Para 3
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (2,1) $(75.1\%,\ 3.2\times 10^{-2})$ (7,2) $(66.8\%,\ 3.3\times 10^{-3})$ (8,2) $(71.1\%,\ 4.2\times 10^{-3})$
2 (6,2) $(78.2\%,\ 2.1\times 10^{-2})$ (7,2) $(72.8\%,\ 5.4\times 10^{-4})$ (8,2) $(79.6\%,\ 1.0\times 10^{-3})$
3 (6,2) $(79.1\%,\ 9.5\times 10^{-3})$ (7,2) $(77.2\%,\ 9.0\times 10^{-4})$ (8,2) $(84.3\%,\ 2.1\times 10^{-4})$
4 (6,2) $(78.7\%,\ 2.4\times 10^{-3})$ (7,2) $(80.1\%,\ 5.7\times 10^{-5})$ (8,2) $(86.4\%,\ 1.0\times 10^{-4})$
5 (6,2) (77.7%, $7\%,\ 2.9\times 10^{-5}) (7$,2) $(82.0\%,\ 5.8\times 10^{-5})$ (8,2) $(87.2\%,\ 9.5\times 10^{-5})$
6 (6,2) $(76.7\%,\ 5.4\times 10^{-5})$ (7,2) $(83.0\%,\ 6.1\times 10^{-5})$ (8,2) $(87.5\%,\ 9.5\times 10^{-5})$
7 (6,2) $(76.0\%,\ 3.8\times 10^{-5})$ (7,2) $(83.6\%,\ 6.3\times 10^{-5})$ (8,2) $(87.6\%,\ 9.6\times 10^{-5})$
8 (6,2) $(75.4\%,\ 3.7\times 10^{-5})$ (7,2) $(83.9\%,\ 6.4\times 10^{-5})$ (8,2) $(87.7\%,\ 9.6\times 10^{-5})$
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.40$.
(d) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$, with $P=0.40$.
Figure 4: Performance for ESPD with initialized parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$ ((a), (b)) and $P=0.40$ ((c), (d)), respectively.
3.1.3 Choice of initialized SPDs
Notably, the requirements imposed on the SPD used for initialization are minimal, which further lowers the practical barrier to implementation. As evidenced by the comparison between Tables 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the initialized SPD, provided that it lies within a certain convergence interval. The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point. Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$. Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient. Indeed, numerical simulations suggest that both SPDs with low DE and low DCR[29], as well as those with moderate DE and relatively high DCR[28], can serve as viable initializations.
In summary, the ESPD scheme requires only a conventional room-temperature SPD for initialization, which is routinely achievable with existing technology. For instance, the SPD developed two decades ago [28] is sufficient.
3.1.4 Auxiliary Sources
Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, implementing the ESPD scheme may not require single photon sources. The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, regardless of the exact number of photons or the presence of unwanted $|0\rangle$. Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required $|1\rangle$ [^3] Since the employment of the filter block, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects. Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary paths, which will be filtered out and thereby have no effect even if the source is multiple-photon.. However, a rigorous formulation of this effect, including its impact on the controlled operations, remains an open problem and is left for future investigation.
3.2 Limitations and Future Works
Despite the significance of the ESPD framework, some limitations remain and warrant future investigation.
3.2.1 Theoretical Completeness
The theoretical framework established in Section 2.2 remains an area for deeper exploration. While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required. Given the inherent mathematical difficulty of this non-linear and time-variant system, and the fact that numerical designs already suffice for practical implementations, such specialized theoretical investigations are deferred to future work.
3.2.2 Resource Overhead vs. Environmental Simplicity
While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels. Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections. Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD. For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (in the absence of techniques such as multiplexing), resulting in hundreds of integrated components. While not posing a fundamental physical barrier (as discussed in Section 3.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits. Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57]. Notably, very-large-scale integration of quantum photonic components has already been experimentally demonstrated, with over 2,500 integrated on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes. Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms. For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation. The SNSPD paradigm mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K). The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration. By leveraging large-scale photonic platforms, thousands of components can be lithographically fabricated on a single chip, effectively converting high operational costs and environmental constraints into a one-time design and fabrication cost. This trade-off is particularly advantageous for applications where size, weight, and power constraints prohibit the use of cryogenics, such as satellite-based quantum communication payloads, mobile quantum nodes, or research laboratories lacking dedicated low-temperature facilities.
3.2.3 Optimization Design
Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section 2.4
Para 4 Para 5 Para 6
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (8,2) $(95.0\%,\ 4.2\times 10^{-3})$ (4,2) $(86.1\%,\ 1.2\times 10^{-3})$ (3,1) $(95.8\%,\ 4.3\times 10^{-2})$
2 (8,4) $(93.5\%,\ 1.4\times 10^{-7})$ (8,4) $(92.6\%,\ 6.4\times 10^{-9})$ (6,4) $(93.9\%,\ 1.2\times 10^{-4})$
3 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.3\%,\ 8.2\times 10^{-10})$ (8,4) $(93.4\%,\ 1.2\times 10^{-9})$
4 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.3\%,\ 8.4\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
5 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
6 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
7 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
8 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Table.2.
The inherent exponential resource scaling discussed in Section 3.2.2 necessitates optimization for future practical deployments. Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$). Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence. As demonstrated in Section 2.4, the choice of design parameters, namely $n$ (the number of auxiliary signals per level) and $k$ (the decision threshold), plays a critical role: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost. Furthermore, as shown in Table 7 and visualized in Fig.5, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
The optimization problem also encompasses the optimal circuit topology. To mitigate cumulative error propagation, minimizing the maximal circuit length is essential. As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1. However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
3.2.4 Physical Integrations and Experiments
Despite the convincing evidence provided by the numerical simulations in Section 2.4, the physical implementation of the ESPD scheme still requires dedicated experimental validation. Reassuringly, as established in Section 3.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization. However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components. Furthermore, the seamless integration of the enhanced detector into various quantum optical tasks (such as QKD) needs full investigation. Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
3.2.5 Count Rate and Timing Jitter
While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors. Specifically, the system’s timing jitter and maximum count rate remain constrained by the intrinsic performance of the initial SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices). However, the ESPD architecture does not introduce significant additional penalties in principle: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, ensuring negligible added temporal broadening. Furthermore, as the final SPDs operate in parallel, the effective system throughput remains comparable to that of the constituent devices. Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics.
4 Conclusion
In summary, the enhanced single-photon detection (ESPD) framework is established to reformulate photon detection as an iteratively enhanced quantum-information-processing task composed of state preparation, controlled operations, projective measurements, and multi-copy decision analysis, rather than as a fixed device-level operation. Within this system-level theoretical framework, detection performance is progressively improved through structured quantum processing using exclusively non-superconducting, room-temperature components.
Numerical simulations, grounded in experimentally reported parameters and commercially available photonic components, demonstrate that the ESPD framework can substantially enhance the effective detection efficiency (DE) (from 59% to 93%) while strongly suppressing the effective dark count rate (DCR) (from $10^{-2}$ to $10^{-9}$) even when initialized with a conventional room-temperature SPD reported two decades ago. The resulting effective performance metrics are comparable to those of state-of-the-art superconducting single-photon detectors, significantly relaxing the minimal tolerable channel transmission rate in quantum communication applications.
Beyond performance enhancement, the ESPD framework makes explicit the fundamental trade-off between resource overhead and operating conditions, replacing cryogenic infrastructure requirements with system-level integration complexity. Although the realization of large-scale ESPD architectures would require further advances in device integration and experimental implementation, the theoretical formulation and parameter studies presented here establish a concrete foundation for the framework and provide concrete guidance for future experimental efforts.
In essence, this work establishes a conceptually complete and theoretically grounded framework for high-performance single-photon detection at room temperature. It demonstrates that intrinsic device-level limitations need not be overcome solely through component replacement, but can instead be effectively compensated by system-level architectural design. As such, the ESPD framework offers a viable pathway toward next-generation photon detection and provides a general methodology for rethinking performance limits in quantum technologies under practical constraints.
Acknowledgements
The author acknowledges the use of Gemini and ChatGPT (accessed December 2025) to, and only to, support language polishing, flow paragraphs, and correct grammar, to improve readability. All AI-generated suggestions were reviewed, revised, and approved by the authors, who take full responsibility for the accuracy and integrity of the work. All insight, results, methods, analysis, and experiments are ordinarily created and owned by the author.
Code Availability
All data/codes generated or used during this study are included in this published article or can be found at https://github.com/Hao-B-Shu/ESPD.